# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ IX.2 Finite $T_{0}$ errors
Figure IX.9: Difference in R between different $T_{0}$ values on last u ray
In figure IX.9, we show the difference in $R$ between the three pairs of successive $T_{0}$ runs. We see in the graph that this difference decreases by a factor of approximately 10 between two such successive pairs. Notice that $e^{2\kappa_{+}\Delta T_{0}}\approx 10.4$ where $\Delta T_{0}=5$ is the difference in $T_{0}$ between two successive runs, which nicely agrees with this factor of $\sim 10$.
@@
Figure IX.10: Difference in S between different $T_{0}$ values on last u ray
In figure IX.10, we show the difference in $S$ between the three pairs of successive $T_{0}$ runs. We can see a similar behavior for $S$ as it was for $R$: the gaps here again correspond to the error being reduced by a factor of approximately 10. Here the error in $T_{0}=-30$ is approximately of order $10^{-5}$.
Figure IX.11: Difference in $R_{,v}$ between different $T_{0}$ values on last u ray
In the following Figure IX.11, we see the differences in $R_{,v}$ between pairs of successive $T_{0}$ runs. In the region $v>600$ we see that the round-off error starts to dominate over the finite-$T_{0}$ error. The error in $T_{0}=-30$ is found to be of order $10^{-10}$ (or smaller) in region 3.
Figure IX.12: Difference in $S_{,v}$ between different $T_{0}$ values on last u ray
Moving on to Figure IX.12, which shows the corresponding differences in the values of $S_{,v}$, we again see that the roundoff error dominates at $v>600$. The finite-$T_{0}$ error in $T_{0}=-30$ is of order $10^{-7}$.
@@
X Independence from the detail of sources
As was already mentioned in subsection IV.8, our analytical approximation for region 3 only depends on the four input functions of $M(v)$, $Q(v)$, $\widetilde{T}_{vv}^{(-)}(v)$ and $\widetilde{T}_{uu}^{(-)}(v)$. Once $M(v)$ and $Q(v)$ are prescribed, they directly determine $\widetilde{T}_{vv}^{(+)}$, and $\widetilde{T}_{uu}^{(+)}=0$ from regularity. Then $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ are in principle determined by the evolution from the EH to the near-IH region (with $\widetilde{T}_{vv}^{(+)}$ and $\widetilde{T}_{uu}^{(+)}=0$ serving as initial values) – in a manner that depends on the sources $Z_{R}$ and $Z_{S}$. Thus, these sources do affect $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ – but otherwise the analytical approximation is independent of these sources. Our goal in this section is to numerically check this independence from the details of sources.
To check this independence, we run the numerical code with three different sets of sources which all lead to (approximately) the same values of $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$, and we show that the results for $R_{,v}$$R_{,u}$$S_{,v}$$S_{,u}$ are the same in all three cases – meaning they are independent of the specific choice of the sources. We use the parametrization given in Eqs. (III.23,III.24) for $Z_{R}$ and $Z_{S}$ in terms of the four parameters $Z_{R0}$, $Z_{Rr}$, $Z_{S0}$, $Z_{Sr}$.
In all three runs, we choose $\alpha=1$ for $\delta Q$, and $\frac{Q}{M}=0.8$. The three sets of source parameters used in these three runs are given in Table 1.
@@ X Independence from the detail of sources
We found empirically that what matters for the final values of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ in region 3 is the combination $Z_{S0}+Z_{Sr}-\frac{1}{2}Z_{Rr}$, as displayed at the last column. (This empirical expression may be explored analytically, but this is beyond the scope of this research.)
In the first stage we shall check that in region 3 we have the same $\widetilde{T}_{uu}$ and $\widetilde{T}_{vv}$ for all three sets of parameters.
Figure X.1: $\widetilde{T}_{uu}$ for the three different runs on the last u ray
Figure X.1 shows that $\widetilde{T}_{uu}$ is approximately equal for the different runs (the curves overlap).
Figure X.2: $\widetilde{T}_{uu}$ for the three different runs, in region 3, on the last u ray
In the next Figure X.2 we zoom in to region 3 where we compare $\widetilde{T}_{uu}$ for the three different runs, and we can see they are pretty close to each other, with difference of order $\sim 10^{-5}$, to be compared with the $\sim 10^{-3}$ value of $\widetilde{T}_{uu}$.
Figure X.3: $\widetilde{T}_{vv}$ for the three different runs on the last u ray
In the following Figure X.3, we show that $\widetilde{T}_{vv}$ is also approximately equal.
Figure X.4: $S_{,v}$ for the three different runs, compared to $\kappa_{-}(v)$ on the last u ray
We move on to show $S_{,v}$ along with its expected value $\kappa_{-}(v)$ in Figure X.4. The analytical approximation works well for all runs.
Figure X.5: $S_{,v}$ for the three different runs, compared to $\kappa_{-}(v)$, zoomed in on the last u ray
We zoom in to the region $800\leq v\leq 900$ in Figure X.5 to show how close are the three different runs.
Figure X.6: $S_{,u}$ for the three different runs, compared to $\kappa_{-}^{u}$ on the last v ray
We continue to show $S_{,u}$ and its expected value $\kappa_{-}^{u}(u)$ in Figure X.6. All the three equivalent runs are close together in region 3, near their expected value of $\kappa_{-}^{u}(u)$.
Figure X.7: $R_{,v}$ for the three different runs, compared to $\frac{\widetilde{T}_{vv}}{\kappa_{-}^{v}}$ on the last u ray
We turn now to explore the derivatives of $R$. Figure X.7 shows both $R_{,v}$ and its analytical approximation $\frac{\widetilde{T}_{vv}}{\kappa_{-}^{v}}$ for each run. All the three runs and their approximations are close together.
Figure X.8: $R_{,u}$ for the three different runs, compared to $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ on the last v ray
Subsequently, we show in Figure X.8 both $R_{,u}$ and its analytical approximation $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ for each run. All curves are close together, showing that the approximation is good.
@@
XI Three additional runs with different sources
We shall now extend the validity test of the analytical approximation to three additional sets of source parameters – this time we chose parameter sets which lead to different values of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ in region 3 (unlike what we did in section X). Table 2 details the various source parameters ($Z_{R0}$, $Z_{Rr}$, $Z_{S0}$, $Z_{Sr}$$\alpha$) and $\frac{Q}{M}$ used in each run. Run0 was detailed in section VII. Run1 was detailed in section VIII. Our focus here is runs 2 through 4.
In all three runs we shall display $R(v)$,$R(u)$, and then the fluxes $\widetilde{T}_{vv}(v)$ and $\widetilde{T}_{uu}(u)$, followed by the 4 derivatives $S_{,v}(v)$$S_{,u}(u)$ and $R_{,v}(v)$$R_{,u}(u)$. All these quantities will be displayed along either the last $u=const$ or $v=const$ rays depending on their arguments (being $u$ or $v$).
As we shall see in all these 3 runs, the analytical approximation works well.
@@ XI.1 Run2
We can see in Figure XI.2a that $\widetilde{T}_{vv}>0$ in region 3, and from Figure XI.2b $\widetilde{T}_{uu}<0$ in region 3. This explains why in Figure XI.1a $R$ is decreasing with $v$ ($R_{,v}<0$) in region 3, while in Figure XI.1b $R$ is increasing with $u$ ($R_{,u}>0$) in region 3.
Figure XI.3a shows the value of $S_{,v}$ is close to its analytical approximation. Similarly, in Figure XI.3b we can see that $S_{,u}$ is also close to $\kappa_{-}^{u}$, as expected from the analytical approximation.
Figure XI.4a compares the value of $R_{,v}$ to the analytical approximation, and demonstrates that they are close. Figure XI.4b compares the value of $R_{,u}$ to the analytical approximation, showing they are close too.
Figure XI.1: R for run2
@@
Figure XI.2: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ for run2
@@
Figure XI.3: $S_{,v}$ and $S_{,u}$ and their analytical approximation for run2
@@
Figure XI.4: $R_{,v}$ and $R_{,u}$ and their analytical approximation for run2
@@ XI.2 Run3
As can be seen in Figure XI.5a, $R$ reaches its minimum value at the end of region 2 and then drifts upwards in region 3. This minimal value of $R=r_{-}^{2}=0.04$ is quite small (compared to e.g. in Figure XI.1a) for a simple reason: Because $\frac{Q}{M}$ is now smaller, $r_{-}$ is smaller than in previous runs when $\frac{Q}{M}$ was chosen to be 0.8.
Figures XI.6a and XI.6b respectively show that both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are <0<0. Because of that, $R_{,u}>0$ and also $R_{,v}>0$ in region 3, as can be seen by the increasing $R$ in Figures XI.5a, XI.5b.
Figures XI.7,XI.8 show the quantities $S_{,v}$$S_{,u}$ and $R_{,v}$$R_{,u}$ respectively, compared with their analytical approximation values. It can be seen that the analytical approximation is close to the numerically computed values for all these quantities.
Figure XI.5: R for run3
@@
Figure XI.6: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ for run3
@@
Figure XI.7: $S_{,v}$ and $S_{,u}$ and their analytical approximation for run 3
@@
Figure XI.8: $R_{,v}$ and $R_{,u}$ and their analytical approximation for run3
@@
XI.3 Run4
This run is done with source parameters taken to be $Z_{R0}=Z_{S0}$=1=1, $Z_{Rr}$$Z_{Sr}=$$\frac{1}{2}$, and $\alpha=-\frac{1}{2}$ for $\delta Q$ and with $\frac{Q}{M}=0.8$. In this run, unlike all other runs, the maximum $v$ is taken to be 482 (compared with 900 in all previous runs) to avoid (or at least reduce) complications that may arise from the singularity that is going to form in this case (see below) – except for Figure XI.10b where $v$ is taken to be 600 in order to show where the singularity forms on this null ray.
Figures XI.11a and XI.11b respectively show that this time, both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are >0>0 in region 3. Because of that, both $R_{,v}$ and $R_{,u}$ are negative there. Correspondingly, Figure XI.9 shows that, at u=60u=-60, $R$ is decreasing and reaching zero at $v=482$. Likewise, Figure XI.10b shows, at $v=600$, that $R$ is also decreasing as a function of $u$, reaching zero at u=274u=-274, which is marked by a short red line. This vanishing of $R$ – seen in all these figures – expresses the formation of a spacelike $R=0$ singularity, which we shall further discuss at the end of this subsection. We also show Figure XI.10a, which shows that at $v=482$, $R$ is decreasing all the way to zero at u=60u=-60. We chose to show quantities on this ray of $v=482$ for the rest of this case to avoid crossing the singularity, as this is the value of $v$ at which $R$ vanishes along the last-$u$ ray u=60u=-60.
Figure XI.12a and XI.12b respectively show $S_{,v}$ and $S_{,u}$ versus their analytical approximations $\kappa_{-}(v)$ and $\kappa_{-}^{u}(u)$, again showing the good quality of the analytical approximation.
Figure XI.13a shows $R_{,v}$ compared with its analytical approximation, and they are close but have a small gap which is this time bigger relative to the previous runs. A similar gap is also seen in Figure XI.13b which shows $R_{,u}$ compared with its analytical approximation. We shall now explore this gap, and show that this deviation from the analytical approximation is a finite-$\varepsilon_{0}$ deviation (which in this specific case happens to be bigger than in the previous cases).
Figure XI.9: $R(v)$ for run4 on last u ray
Figure XI.10: $R(u)$ for run4 along two $v=const$ rays
@@ XI.3 Run4
Figure XI.11: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ for run4
@@
Figure XI.12: $S_{,v}$ and $S_{,u}$ and their analytical approximation for run4
@@
Figure XI.13: $R_{,v}$ and $R_{,u}$ and their analytical approximation for run4
@@
To this end, we shall run the same case with $\varepsilon_{0}=\frac{1}{500}$ as well, and compare the deviations from the analytical approximation obtained in both values of $\varepsilon_{0}$, namely $\frac{1}{500}$ and $\frac{1}{1000}$, similar to subsection VII.4. Figure XI.14 presents this deviation in $R_{,v}$ for these two values of $\varepsilon_{0}$, with the $\varepsilon_{0}=\frac{1}{500}$ case scaled by $\frac{1}{4}$. The blue and orange curves overlap, demonstrating that this deviation from the analytical approximation scales as $\varepsilon_{0}^{2}$ – indicating that this is indeed a finite-$\varepsilon_{0}$ deviation.
Figure XI.14: $R_{,v}+\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ with two $\varepsilon_{0}$ and scaled by $\frac{1}{4}$ in run4 on last u ray
Figure XI.15: Spacetime diagram for $R=0$ singularity
In this case (and unlike the previous cases we explored above), because both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are positive, a spacelike $R=0$ singularity forms at the middle of the numerical domain, blocking access to the would-be CH. The spacetime diagram corresponding to this situation is shown in Figure XI.15.777There is a possibility that the two fluxes will have opposite signs in region 3. In this case, as it turns out, what matters is the sign of $\widetilde{T}_{uu}$: if it is negative (as it is e.g. in run 2 above), then a CH will form, and if it is positive, a spacelike $R=0$ singularity will form instead.
XII Conclusion

### v2

@@ 9.2 Finite $T_{0}$ errors
Figure 9.9: Difference in R between different $T_{0}$ values on last u ray
In figure 9.9, we show the difference in $R$ between the three pairs of successive $T_{0}$ runs. We see in the graph that this difference decreases by a factor of approximately 10 between two such successive pairs. Notice that $e^{2\kappa_{+}\Delta T_{0}}\approx 10.4$ where $\Delta T_{0}=5$ is the difference in $T_{0}$ between two successive runs, which nicely agrees with this factor of $\sim 10$.
@@
Figure 9.10: Difference in S between different $T_{0}$ values on last u ray
In figure 9.10, we show the difference in $S$ between the three pairs of successive $T_{0}$ runs. We can see a similar behavior for $S$ as it was for $R$: the gaps here again correspond to the error being reduced by a factor of approximately 10. Here the error in $T_{0}=-30$ is approximately of order $10^{-5}$.
Figure 9.11: Difference in $R_{,v}$ between different $T_{0}$ values on last u ray
In the following Figure 9.11, we see the differences in $R_{,v}$ between pairs of successive $T_{0}$ runs. In the region $v>600$ we see that the round-off error starts to dominate over the finite-$T_{0}$ error. The error in $T_{0}=-30$ is found to be of order $10^{-10}$ (or smaller) in region 3.
Figure 9.12: Difference in $S_{,v}$ between different $T_{0}$ values on last u ray
Moving on to Figure 9.12, which shows the corresponding differences in the values of $S_{,v}$, we again see that the roundoff error dominates at $v>600$. The finite-$T_{0}$ error in $T_{0}=-30$ is of order $10^{-7}$.
@@
10 Independence from the detail of sources
As was already mentioned in subsection 4.8, our analytical approximation for region 3 only depends on the four input functions of $M(v)$, $Q(v)$, $\widetilde{T}_{vv}^{(-)}(v)$ and $\widetilde{T}_{uu}^{(-)}(v)$. Once $M(v)$ and $Q(v)$ are prescribed, they directly determine $\widetilde{T}_{vv}^{(+)}$, and $\widetilde{T}_{uu}^{(+)}=0$ from regularity. Then $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ are in principle determined by the evolution from the EH to the near-IH region (with $\widetilde{T}_{vv}^{(+)}$ and $\widetilde{T}_{uu}^{(+)}=0$ serving as initial values) – in a manner that depends on the sources $Z_{R}$ and $Z_{S}$. Thus, these sources do affect $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ – but otherwise the analytical approximation is independent of these sources. Our goal in this section is to numerically check this independence from the details of sources.
To check this independence, we run the numerical code with three different sets of sources which all lead to (approximately) the same values of $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$, and we show that the results for $R_{,v}$$R_{,u}$$S_{,v}$$S_{,u}$ are the same in all three cases – meaning they are independent of the specific choice of the sources. We use the parametrization given in Eqs. (3.23,3.24) for $Z_{R}$ and $Z_{S}$ in terms of the four parameters $Z_{R0}$, $Z_{Rr}$, $Z_{S0}$, $Z_{Sr}$. In all three runs, we choose $\alpha=1$ for $\delta Q$, and $\frac{Q}{M}=0.8$. The three sets of source parameters used in these three runs are given in Table 1.
@@ 10 Independence from the detail of sources
We found empirically that what matters for the final values of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ in region 3 is the combination $Z_{S0}+Z_{Sr}-\frac{1}{2}Z_{Rr}$, as displayed at the last column. (This empirical expression may be explored analytically, but this is beyond the scope of this research.) In the first stage we shall check that in region 3 we have the same $\widetilde{T}_{uu}$ and $\widetilde{T}_{vv}$ for all three sets of parameters.
Figure 10.1: $\widetilde{T}_{uu}$ for the three different runs on the last u ray
Figure 10.1 shows that $\widetilde{T}_{uu}$ is approximately equal for the different runs (the curves overlap).
Figure 10.2: $\widetilde{T}_{uu}$ for the three different runs, in region 3, on the last u ray
In the next Figure 10.2 we zoom in to region 3 where we compare $\widetilde{T}_{uu}$ for the three different runs, and we can see they are pretty close to each other, with difference of order $\sim 10^{-5}$, to be compared with the $\sim 10^{-3}$ value of $\widetilde{T}_{uu}$.
Figure 10.3: $\widetilde{T}_{vv}$ for the three different runs on the last u ray
In the following Figure 10.3, we show that $\widetilde{T}_{vv}$ is also approximately equal.
Figure 10.4: $S_{,v}$ for the three different runs, compared to $\kappa_{-}(v)$ on the last u ray
We move on to show $S_{,v}$ along with its expected value $\kappa_{-}(v)$ in Figure 10.4. The analytical approximation works well for all runs.
Figure 10.5: $S_{,v}$ for the three different runs, compared to $\kappa_{-}(v)$, zoomed in on the last u ray
We zoom in to the region $800\leq v\leq 900$ in Figure 10.5 to show how close are the three different runs.
Figure 10.6: $S_{,u}$ for the three different runs, compared to $\kappa_{-}^{u}$ on the last v ray
We continue to show $S_{,u}$ and its expected value $\kappa_{-}^{u}(u)$ in Figure 10.6. All the three equivalent runs are close together in region 3, near their expected value of $\kappa_{-}^{u}(u)$.
Figure 10.7: $R_{,v}$ for the three different runs, compared to $\frac{\widetilde{T}_{vv}}{\kappa_{-}^{v}}$ on the last u ray
We turn now to explore the derivatives of $R$. Figure 10.7 shows both $R_{,v}$ and its analytical approximation $\frac{\widetilde{T}_{vv}}{\kappa_{-}^{v}}$ for each run. All the three runs and their approximations are close together.
Figure 10.8: $R_{,u}$ for the three different runs, compared to $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ on the last v ray
Subsequently, we show in Figure 10.8 both $R_{,u}$ and its analytical approximation $\frac{\widetilde{T}_{uu}}{\kappa_{-}^{u}}$ for each run. All curves are close together, showing that the approximation is good.
@@
11 Three additional runs with different sources
We shall now extend the validity test of the analytical approximation to three additional sets of source parameters – this time we chose parameter sets which lead to different values of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ in region 3 (unlike what we did in section 10). Table 2 details the various source parameters ($Z_{R0}$, $Z_{Rr}$, $Z_{S0}$, $Z_{Sr}$$\alpha$) and $\frac{Q}{M}$ used in each run. Run0 was detailed in section 7. Run1 was detailed in section 8. Our focus here is runs 2 through 4.
In all three runs we shall display $R(v)$,$R(u)$, and then the fluxes $\widetilde{T}_{vv}(v)$ and $\widetilde{T}_{uu}(u)$, followed by the 4 derivatives $S_{,v}(v)$$S_{,u}(u)$ and $R_{,v}(v)$$R_{,u}(u)$. All these quantities will be displayed along either the last $u=const$ or $v=const$ rays depending on their arguments ($u$ or $v$). As we shall see in all these 3 runs, the analytical approximation works well.
@@ 11.1 Run2
We can see in Figure 11.2a that $\widetilde{T}_{vv}>0$ in region 3, and from Figure 11.2b $\widetilde{T}_{uu}<0$ in region 3. This explains why in Figure 11.1a $R$ is decreasing with $v$ ($R_{,v}<0$) in region 3, while in Figure 11.1b $R$ is increasing with $u$ ($R_{,u}>0$) in region 3. Figure 11.3a shows the value of $S_{,v}$ is close to its analytical approximation. Similarly, in Figure 11.3b we can see that $S_{,u}$ is also close to $\kappa_{-}^{u}$, as expected from the analytical approximation. Figure 11.4a compares the value of $R_{,v}$ to the analytical approximation, and demonstrates that they are close. Figure 11.4b compares the value of $R_{,u}$ to the analytical approximation, showing they are close too.
Figure 11.1: R for run2
@@
Figure 11.2: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ for run2
@@
Figure 11.3: $S_{,v}$ and $S_{,u}$ and their analytical approximation for run2
@@
Figure 11.4: $R_{,v}$ and $R_{,u}$ and their analytical approximation for run2
@@ 11.2 Run3
As can be seen in Figure 11.5a, $R$ reaches its minimum value at the end of region 2 and then drifts upwards in region 3. This minimal value of $R=r_{-}^{2}=0.04$ is quite small (compared to e.g. in Figure 11.1a) for a simple reason: Because $\frac{Q}{M}$ is now smaller, $r_{-}$ is smaller than in previous runs when $\frac{Q}{M}$ was chosen to be 0.8. Figures 11.6a and 11.6b respectively show that both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are <0<0. Because of that, $R_{,u}>0$ and also $R_{,v}>0$ in region 3, as can be seen by the increasing $R$ in Figures 11.5a, 11.5b.Figures 11.7,11.8 show the quantities $S_{,v}$$S_{,u}$ and $R_{,v}$$R_{,u}$ respectively, compared with their analytical approximation values. It can be seen that the analytical approximation is close to the numerically computed values for all these quantities.
Figure 11.5: R for run3
@@
Figure 11.6: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ for run3
@@
Figure 11.7: $S_{,v}$ and $S_{,u}$ and their analytical approximation for run 3
@@
Figure 11.8: $R_{,v}$ and $R_{,u}$ and their analytical approximation for run3
@@
11.3 Run4
This run is done with source parameters taken to be $Z_{R0}=Z_{S0}$=1=1, $Z_{Rr}$$Z_{Sr}=$$\frac{1}{2}$, and $\alpha=-\frac{1}{2}$ for $\delta Q$ and with $\frac{Q}{M}=0.8$. In this run, unlike all other runs, the maximum $v$ is taken to be 482 (compared with 900 in all previous runs) to avoid (or at least reduce) complications that may arise from the singularity that is going to form in this case (see below) – except for Figure 11.10b where $v$ is taken to be 600 in order to show where the singularity forms on this null ray.
Figures 11.11a and 11.11b respectively show that this time, both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are >0>0 in region 3. Because of that, both $R_{,v}$ and $R_{,u}$ are negative there. Correspondingly, Figure 11.9 shows that, at u=60u=-60, $R$ is decreasing and reaching zero at $v=482$. Likewise, Figure 11.10b shows, at $v=600$, that $R$ is also decreasing as a function of $u$, reaching zero at u=274u=-274, which is marked by a short red line. This vanishing of $R$ – seen in all these figures – expresses the formation of a spacelike $R=0$ singularity, which we shall further discuss at the end of this subsection. We also show Figure 11.10a, which shows that at $v=482$, $R$ is decreasing all the way to zero at u=60u=-60. We chose to show quantities on this ray of $v=482$ for the rest of this case to avoid crossing the singularity, as this is the value of $v$ at which $R$ vanishes along the last-$u$ ray u=60u=-60. Figure 11.12a and 11.12b respectively show $S_{,v}$ and $S_{,u}$ versus their analytical approximations $\kappa_{-}(v)$ and $\kappa_{-}^{u}(u)$, again showing the good quality of the analytical approximation. Figure 11.13a shows $R_{,v}$ compared with its analytical approximation, and they are close but have a small gap which is this time bigger relative to the previous runs. A similar gap is also seen in Figure 11.13b which shows $R_{,u}$ compared with its analytical approximation. We shall now explore this gap, and show that this deviation from the analytical approximation is a finite-$\varepsilon_{0}$ deviation (which in this specific case happens to be bigger than in the previous cases).
Figure 11.9: $R(v)$ for run4 on last u ray
Figure 11.10: $R(u)$ for run4 along two $v=const$ rays
@@ 11.3 Run4
Figure 11.11: $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ for run4
@@
Figure 11.12: $S_{,v}$ and $S_{,u}$ and their analytical approximation for run4
@@
Figure 11.13: $R_{,v}$ and $R_{,u}$ and their analytical approximation for run4
@@
To this end, we shall run the same case with $\varepsilon_{0}=\frac{1}{500}$ as well, and compare the deviations from the analytical approximation obtained in both values of $\varepsilon_{0}$, namely $\frac{1}{500}$ and $\frac{1}{1000}$, similar to subsection 7.4. Figure 11.14 presents this deviation in $R_{,v}$ for these two values of $\varepsilon_{0}$, with the $\varepsilon_{0}=\frac{1}{500}$ case scaled by $\frac{1}{4}$. The blue and orange curves overlap, demonstrating that this deviation from the analytical approximation scales as $\varepsilon_{0}^{2}$ – indicating that this is indeed a finite-$\varepsilon_{0}$ deviation.
Figure 11.14: $R_{,v}+\frac{\widetilde{T}_{vv}}{\kappa_{-}}$ with two $\varepsilon_{0}$ and scaled by $\frac{1}{4}$ in run4 on last u ray
Figure 11.15: Spacetime diagram for $R=0$ singularity
In this case (and unlike the previous cases we explored above), because both $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$ are positive, a spacelike $R=0$ singularity forms at the middle of the numerical domain, blocking access to the would-be CH. The spacetime diagram corresponding to this situation is shown in Figure 11.15.This $r=0$ spacelike singularity that forms inside the semiclassical charged black hole resembles the spacelike singularity found in Ref [6], although it must be recalled that they studied classical perturbations whereas here we consider the backreaction of a quantum field.  [^8] There is a possibility that the two fluxes will have opposite signs in region 3. In this case, as it turns out, what matters is the sign of $\widetilde{T}_{uu}$: if it is negative (as it is e.g. in run 2 above), then a CH will form, and if it is positive, a spacelike $R=0$ singularity will form instead.
12 Conclusion